40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 204 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  7078   Thu Aug 2 11:09:52 2012 EricSummaryLSCFree-Swinging Michelson Measurements

To take the free swinging Michelson measurements for the interferometer calibration Jamie aligned the beam splitter with ITMX and ITMY. I recorded the GPS time (1027827100 and for several hundred seconds later) when the Michelson was aligned in order to look at the correct data. I then copied the python script nds-test.py from Jamie, and modified it to take and plot data from C1:LSC-AS55_Q_ERR_DQ offline. I used dataviewer to verify that C1:LSC-AS55_Q_ERR_DQ and C1:LSC-AS55_Q_ERR were recording the same signal, and to check that I was taking the correct data with NDS. Taking data online worked as well, but it was easier to use a time when the Michelson was known to be free-swinging and take data offline. Attached is some sample data while free-swinging, with time in GPS time.

Attachment 1: free_swing_MICH.png
free_swing_MICH.png
  7077   Thu Aug 2 04:58:00 2012 MashaUpdatePEM70 Meter Long Guralp 1 Cable

The parts Jenne and I ordered arrived today, so we made a long cable for Guralp 1 using a 24 + 1 wire 70 meter long cable, a female 37-pin DSub, and a 26-pin milspec. The pin map is the same as the one I specified in my previous E-log. I soldered both the milspec attachment and the DSub attachment, and used a Multimeter to check the connectivity of the cables. 20 of 20 connections worked (beeped), so I plugged  the cable into the Gurlap 1 seismometer and the Guralp box.

The time series comparison for the two cables

Old cable:

BeforeCable.png

New cable: (I had to move GUR 1, so it's still stabilizing in the X and Y time series)

 

 AfterCable.pngNew

The current signal spectrum

 

 NewCableFreq.png

The BLRMS on the seismic strip also look similar using the two cables - it's more visible on the wall, but I will include a StripTool picture:

New Cable BLRMS (similar to old cable BLRMS)

 NewCableStrip.png

  7076   Thu Aug 2 03:06:57 2012 SashaUpdateSimulationsLS Plant (LSP) is officially ONLINE

My ls plant compiled!! The RCG code can now be found in /opt/rtcds/rtscore/tags/advLigoRTS-2.5. I uploaded a copy of c1lsp.mdl onto the svn.

The weird "failed to connect" error was due to the fact that I named my inputs the same thing as my goto/from tags, so the RCG got confused. Once I renamed my inputs, it worked! I'm not sure what happened to the original "not enough parts" error; it didn't appear a single time during the rebuilding process. Anyway, I made the PDH block much neater, though the lines between PDH and ADC are looking wonky (this is purely an aesthetic problem, not a "oh god my simulation will DIE right now if I don't fix it" problem). I'll fix it in the morning; screenshot attached!

The original c1lsp was kind of sad. I updated it extensively and brought it into the modern era with color. The original c1lsp.mdl should also be on the svn. Tommorow, I'll get started on figuring out how to get LIGO specific noises from white noise.

Attachment 1: Screenshot-1.png
Screenshot-1.png
  7075   Thu Aug 2 00:43:55 2012 JenneUpdateASSMajor cleanup of the ASS model

[Jamie, Jenne]

We are trying to figure out what the story is with the ASS, and in order to make it more human parse-able, we cleaned up the c1ass.mdl. 

So far, we have made no changes to how the signals are routed.  The local oscillators from each lockin still get summed, and go directly to the individual optics, and the demodulated signals from each lockin go through the sensing matrix, the DoF filters, then the control output matrix, and then on to the individual optics.  So far, so good.

Much of the cleanup work involved making a big library part, which is then called once for PIT and once for YAW in the ass top level, rather than have 2 code-copies, which give Jamie conniptions.  Inside the library part GoTo and From tags were used, rather than having all the lines cross over one another in a big spaghetti mess.

One of the big actual changes to the ass was some name changes. Rather than having mysterious "ASS-LOCKIN(1-30)", they are now named something like "ASS-PIT_LOCKIN_ETMY_TRY", indicating that this is in the pitch set, actuating on ETMY, and looking at TRY for the demodulated signal.  The "DOF" channels are similar to what they were, although we would like to change them in the future.....more on this potential change later.  Previously they were "ASS-DOF(1-10)", but now they are "ASS-PIT_DOF(1-5)" and "ASS-YAW_DOF(1-5)". This channel naming, while it makes things make more sense, and is easier to understand, means that all of the ASS scripts need to be fixed.  However, they all needed updating / upgrading anyway, so this isn't the end of the world.

This channel name fixing also included updating names of IPC (shmem/dolphin/rfm things) blocks, which required matching changes in the SUS, RFM and LSC models.  All 4 models (ASS, SUS, RFM, LSC) were recompiled and installed.  They all seem fine, except there appears to be a dolphin naming mismatch between OAF and SUS that showed up when the SUS was recompiled, which presumably it hadn't been in a while.  We need to figure this out, but maybe not tonight.  Den, if you have time, it would be cool if you could take a look at the OAF and SUS models to make sure the names match when sending signals back and forth.

______________________

We also had a long chat about the deeper meaning of the ASS. 

What should we be actuating on, and what should we be sensing?  A potential thought is to rename our DOF channels to actual DoF names: input axis translation, input axis angle, cavity axis translation, cavity axis angle.  Then actuate the dither lines on a cavity degree of freedom, sense the influence on TRX, TRY and an LSC PD (as is currently done), then actuate on the cavity degree of freedom.

Right now, it looks like the actuation is for individual optics, the sensing is the influence on TRX, TRY and an LSC PD, then actuate on a cavity degree of freedom.  So the only change with the new idea is that we actuate in the DoF basis, not the optics basis.  So the Lockin local oscillators would go through the control output matrix.  This makes more sense in my head, but Jamie and I wanted to involve more people in the conversation before we commit. 

The next question would be: how do we populate the control output matricies?  Valera (or someone) put something in there, but I couldn't find anything on the elog about how those values were measured/calculated/came-to-be.  Any ideas?  If we want to dither and then push on isolated degrees of freedom, we need to know how much moving of which optics affects each DoF.  Is this something we should do using only geometry, and our knowledge of optic placements and relative distances, or is this measurable?

  7074   Wed Aug 1 22:37:21 2012 JenneUpdateASSFilters installed in the ASS

As part of trying to figure out what is going on with the ASS, I wanted to figure out what filters are installed on which lockins.

Each "DoF"(1-6) has a zpk(0.1,0.0001,1)gain(1000), which is a lowpass with 60dB of gain at DC, and unity gain at high frequencies.

For the lockins, since there are so many, I made a spreadsheet to keep track of them (attached). 

So, what's the point?  The point is, I think that all of the LOCKIN_I filter modules should be the same, with a single low pass filter.  The Q filter banks don't matter, since we don't use those signals, and the signals are grounded inside the model.  The phase of each lockin was / should be tuned such that all of the interesting signal goes to I, and nothing goes to Q.  The SIG filter modules seem okay, in that they're all the same, except for their band pass frequency.  I just need to check to see what frequency the ASS scripts are trying to actuate at, to make sure we're bandpassing the correct things.

 

Attachment 1: ASS_lockin_filters_1Aug2012.pdf
ASS_lockin_filters_1Aug2012.pdf ASS_lockin_filters_1Aug2012.pdf
  7073   Wed Aug 1 18:20:58 2012 JamieUpdateLSCYarm recovered

[Jenne, Jamie]

We recovered lock and alignment of the Y arm.  TRY_OUT is now at ~0.9, after tweaking {I,E}TMY pit/yaw and PZT2.  YARM_GAIN is 0.1.

I saved ITMY, ETMY, and PZT2 alignments in the IFO_ALIGN screen with the new alignment save/restore stuff I got working.

Working on getting Yarm ASS working now...

  7072   Wed Aug 1 17:04:22 2012 JenneUpdateSUSspare SRMU 03 optic moved

Quote:

There was one  NOT MARKED SOS with two broken magnets on its face. This is labeled ???

 While I'm not sure what specific optic this is, I think it's an older optic.  (a) All of the new optics we got from Ramin were enscribed with their #.  (b) This optic appears to have a short arrow scribe line (about the length of the guiderod), and then no scribe line (that I could see through the glass dish) on the other side.  The new optics all have a long arrow scribe line, ~1/2 the full width of the optic, and have clear scribe lines on the opposite side.

  7071   Wed Aug 1 16:38:46 2012 steveUpdateSUSspare SRMU 03 optic moved

SOS optics prepared to be hanged are moved from the South Flow Bench to S15 Clean Cabinet.

SRMU 03 (1-25-2010) specification summery E080460-05-D,  older vintage SRM 01 and PRM 02 (need more specification)

There was one  NOT MARKED SOS with two broken magnets on its face. This is labeled ???

 

This was done to prepare clean space for TIP-TILT drive- test set up.

The existing cable from 1X5 can reach only to the south end:  from whitening filter to satelite amp.  This will be good for future test of suspensions.

We need to make new cable from 1 X 1 to the south end = 40m long

Attachment 1: IMG_1464.JPG
IMG_1464.JPG
Attachment 2: IMG_1465.JPG
IMG_1465.JPG
Attachment 3: IMG_1466.JPG
IMG_1466.JPG
  7070   Wed Aug 1 15:14:20 2012 JenneUpdateIOOPSL Pointing QPD signals lost in late-June 2012

I was looking into why we don't have any light on the PSL pointing QPDs, and it turns out that it has been this way since ~June 29th 2012.  I need to look back in the elog to see what was going on on the PSL table that day, but I suspect it has something to do with Yuta and I, working on the beat setup, since this is all very near that area.

Attached is a plot of the loss of signal on the QPDs.

UPDATE:

 We lost IP POS on the same day as we lost the PSL pointing.  See 2nd attachment.  The _S_Calc is the sum, and it almost looks like the light got near the edge of the diode and just kept falling off until it was gone.  The sum started getting lower on May 16th, and then was gone on June 29th.

So far I've gone back as far as Jan 2012, but I still haven't found any data where we *did* have light on IP ANG.  Sad.

UPDATE, UPDATE (like P.P.S.):  June 29th was the day of the vent...see elog 6895.

Attachment 1: IOO_QPDs_lost_midJune2012.png
IOO_QPDs_lost_midJune2012.png
Attachment 2: IP_QPDs_lost_midJune2012.png
IP_QPDs_lost_midJune2012.png
  7069   Wed Aug 1 15:02:29 2012 JenneUpdateIOOIP POS QPD centered

Jamie went out to look at IP POS, and the beam was *way* off.  Even though our alignment is still rough, we're kind of close right now, so Jamie put the beam back on the QPD, but we need to recenter IPPOS after we get good alignment.

  7068   Wed Aug 1 11:54:59 2012 YaakovSummarySTACISGeophone calibration and open loop gains

This week I've looked into finding an accurate sensitivity for the geophones in the STACIS. I found that when placing a geophone and accelerometer side by side, and using the sensitivity values I had from spec sheets, the readings were very different (see eLog 7054: http://nodus.ligo.caltech.edu:8080/40m/7054).

I cut the shunt resistor off one of the STACIS geos and found it to be 4000 Ohm, which is one of the standard values for this geophone model. When it is connected to the geophone the net resistance is 2000 Ohm (I took a more careful measurement, I took the geophone out). Then the internal coil resistance should be 4000 Ohm, if they are connected in parallel. However, the geophone spec sheet does not have a sensitivity value for this exact scenario, so I'll have to find a different way to determine the calibration (maybe by putting it next to a seismometer with a known sensitivity). So I know for sure that the sensitivity value I was originally using is wrong, because it assumed an internal coil resistance of 380 Ohm, but I have to check if the value I found by forcing the geophones to agree with the accelerometers (eLog 7054 --> 0.25 (m/s)/V) is correct.

I've also been looking again at the open loop gains of the STACIS (see eLog 7058: http://nodus.ligo.caltech.edu:8080/40m/7058). Attached is what TMC, which makes the STACIS, says it should look like (with a 4000 lb load on the STACIS). Today I am taking the open loop gains into higher frequencies to get a better comparison, but the plots look quite similar to what I have so far. So if it is an unstable open loop gain, then it's at least not new.

Attachment 1: 08011201.pdf
08011201.pdf 08011201.pdf
  7067   Wed Aug 1 11:50:49 2012 JamieUpdateCDSadded input monitors to LSC_TRIGGER library part

I added an EPICS monitor to the input of the LSC_TRIGGER part, to allow monitoring the signal used for the trigger.  I then added the monitors to the C1LSC_TRIG_MTRX screen (see below).  This should hopefully aid in setting the trigger levels.

Attachment 1: trigmtrx.png
trigmtrx.png
  7066   Wed Aug 1 11:46:16 2012 MashaSummaryGeneralWeek Summary

 A lot of my time this week was spent struggling with implementing my neural network code in Simulink in order to experiment with using neural network control systems, as Rana suggested. Perhaps I'm inept at S-Functions, but I decided try to use the Model Reference Controller block in Simulink instead of my own code, and experiment with using that to control a driven, damped harmonic oscillator with noise. The block consists of two neural networks, one which is trained on a model of the plant to simulate the plant, and one that it trained on a reference model (which, in my case, is just input -> gain 0 -> output since my desired signal for now is 0. So far, I have managed to adjust the parameters enough to stop the neural network controller from outputting too much force (this causing the amplitude of the oscillator to increase with each iteration), but outputting enough to keep the plant oscillating with a maximal displacement of 2 m/s (with 30 neurons, 100 reference time delays, and 100 plant output time delays). I will continue to work on this, especially with added noise sources, and see how feasible it is to train such a controller to actually perform well.

control.png

20-100-100-Perform.png

As far as classification, I took up that project again, and realized that I have been approaching it the wrong way the whole time - instead of using a neural network to classify an entire time series, I could just classify it online using a recurrent neural network. This, however, meant that my data (which used to be in packets of 900 seconds) had to be parsed through to generate time-dependent desired output vectors. I did this last night, and have been trying various combinations of neuron numbers, time delays, and learning parameters this morning. Below is my current best result for mean square error with time, obtained with 1 neuron in the hidden layer, 50 time delays (so that there are actually 51 neurons feeding into the hidden layer, and a subnetwork of 50 neurons connected to 50 neurons, and learning parameter 0.7). The peak is due to the fact that a large amount of sharp earthquakes occur around that time, essentially giving the neural network a surprise, and causing it to rapidly learn. However, I suspect this sharp rise would decrease if I were to stop decimating my data by a factor of 256, and using all of the inputs as they come in (this, however would be drastically slower). Currently, I have a massive loop running which tries different combinations of neurons and time delays.

1-50-0.7.png

 

In terms of other lab stuff, Jenne and I ordered parts to make a cable for Gurlap-1, and I updated the pin map for Gurlap-1. Also, I wrote my progress report.

 

  7065   Wed Aug 1 10:45:38 2012 SashaSummarySimulationsSURF - Week 6 - Summary

This week, I worked on transferring my Simulink simulation to the RCG. I made all relevant library parts, now under "SASHA library" in the main Simulink library browser. My main concern has been noises - I've added some rudimentry shot noise, amplitude noise, phase noise, and intensity noise. I have yet to add local oscillator noise, and plan to upgrade the existing noises to actually have the PSD they should (using equations from Rana's and Robb Ward's theses). I'm fairly certain this can be achieved by applying the correct transfer function to white noise (a technique I learned from Masha this week!), so the RCG should be able to handle it (theoretically).

I've also been tweaking my main simulation. After a brief (but scary) attempt at adding optical levers, I decided to shelve it in order to focus on noises/RCG simulating. This is not permanent, and I plan to return to them at some point this week or next. My main problem with them was that I knew how to get from optical lever input to pitch/yaw, but had no idea how to get from pitch/yaw to optical lever input. If I had a complete basis for one in terms of the other, I'd be able to, but I don't think this is the way to go. I'm sure there is a good way to do it (it was done SOMEHOW in the original simulation of the suspensions), I just don't know it yet.

In the aftermath of the optical lever semi-disaster, my simulation is once again not really outputting anything, but since it actually worked before adding the optical levers, I'm pretty sure I can get it to work again (this is why its important to use either git or BACKUPS, >.< (or both)).

We also wrote our progress reports this week. Mine includes some discussion on the basics of cavities/the mechanics of the suspensions/brief intro to PDH, and I'll add a section on noises in the next draft. Maybe this'll be of some use to someone someday? One can only hope.

 

 

  7064   Wed Aug 1 10:38:11 2012 EricSummaryGeneralSURF Update

Since my last update I have modified the DARM control loop model to the extent that it resembles the
measured open loop transfer function much more closely
. The phase especially is much more accurate, with
a phase margin of about 35 degrees at the unity gain frequency of 156 Hz. Right now I'm normalizing to
the unity gain frequency still to adjust the gain properly. Using the length response function from the
model, I can calibrate the error signal as well to find the simulated h(t) output. There were a number of
computational problems in calculating the length response function, but I eventually found a work-around.
Attached is an updated plot of the open loop transfer function and the length response function of the model.

This week Jamie showed me around the real-time Simulink models as well. The one specific to my project
is c1cal.mdl
. It takes the output in the form of the error and control signals from c1lsc.mdl as its
input and produces the calibrated signal as output. In order to produce the calibrated signal we need the actuation
function and the inverse of the sensing function for the model as it stands now. We also built, installed,
and restarted the c1cal model because no data was showing up in data viewer, but the problem remained
after this attempt. 

Jamie and I also started on calibrating the interferometer in the traditional way. Jamie aligned the beam
splitter and the input test masses so we could take free-swinging Michelson measurements. However, taking
the data with the nds system appears to be giving different results than what is showing up in data viewer.
The goal of this measurement is to get a value for the peak to peak amplitude of the Michelson error signal.

Attachment 1: OLG_hendries.png
OLG_hendries.png
Attachment 2: Length_response_hendries.png
Length_response_hendries.png
  7063   Wed Aug 1 10:07:16 2012 LizUpdateComputer Scripts / ProgramsWeek 7 Update

Over the past week, I have continued refining the summary pages.  They are now online in their final home, and can be easily accessed from the 40 meter Wiki page!  (It can be accessed by the Daily Summary link under "LOGS").  I have one final section to add plots to (the IFO section is currently still only "dummy" plots) but the rest are showing correct data!  I have many edits to make in order for them to be more intelligible, but they are available for browsing if anyone feels so inclined.

I also spent quite a while formatting the pages so that the days are in PDT time instead of UTC time.  This process was quite time consuming and required modifications in several files, but I tracked my changes with git so they are easy to pinpoint.  I also did a bit of css editing and rewriting of a few html generation functions so that the website is more appealing.  (One example of this is that the graphs on each individual summary page are now full sized instead of a third of the size.

This week, I also worked with the BLRMS mic channels I made.  I edited the band pass and low pass filters that I had created last week and made coherence plots of the channels.  I encountered two major issues while doing this.  Firstly, the coherence of the channels decreases dramatically above 40 Hz.  I will look at this more today, but am wondering why it is the case.  If nothing could be done about this, it would render three of my channels ineffective.  The other issue is that the Nyquist frequency is at 1000 Hz, which is the upper limit of my highest frequency channel (300-1000 Hz).  I am not sure if this really affects the channel, but it looks very different from all of the other channels.  I am also wondering whether the channels below 20 Hz are useful at all, or whether they are just showing noise.

The microphone calibration has been something I have been trying to figure out for quite some time, but I recently found a value on the website that makes the EM172 microphones and has a value for their sensitivity.  I determined the transfer factor from this sensitivity as 39.8107 mV/Pa, although I am not sure if all of the mics will be consistent with this.

  7062   Tue Jul 31 20:59:35 2012 JamieUpdateCDSfixing up MEDM snapshots

I've started to try to fix all the old non-working medm snapshots.  I've made a new snapshot directory, and some new snapshot scripts to handle taking the snapshots, and view old ones.

Snapshots are now in:  /opt/rtcds/caltech/c1/medm/snap

There is one main script: /opt/rtcds/caltech/c1/medm/snap/snapcommands.  This is linked by:

  • updatesnap: update a snapshot
  • viewsnap: view most recent snapshot
  • viewold: view old snapshots

These commands take either the path to the screen in question, or it's relative path to the medm directory (/opt/rtcds/caltech/c1/medm).  The snapshots for a specific screen are stored in a directory specific to the screen, in a place relative to the snap directory that mimics the screens relative path to the overall medm directory.  So for instance, the snap directory for: 

/opt/rtcds/caltech/c1/medm/c1lsc/master/C1LSC_OVERVIEW.adl

is:

controls@rossa:~ 0$ find /opt/rtcds/caltech/c1/medm/snap/c1lsc/master/C1LSC_OVERVIEW/
/opt/rtcds/caltech/c1/medm/snap/c1lsc/master/C1LSC_OVERVIEW/
/opt/rtcds/caltech/c1/medm/snap/c1lsc/master/C1LSC_OVERVIEW/2012-08-01-02:21:34-UTC.png
/opt/rtcds/caltech/c1/medm/snap/c1lsc/master/C1LSC_OVERVIEW/2012-08-01-02:21:04-UTC.png
/opt/rtcds/caltech/c1/medm/snap/c1lsc/master/C1LSC_OVERVIEW/2012-08-01-02:21:23-UTC.png
/opt/rtcds/caltech/c1/medm/snap/c1lsc/master/C1LSC_OVERVIEW/2012-08-01-03:41:34-UTC.png
/opt/rtcds/caltech/c1/medm/snap/c1lsc/master/C1LSC_OVERVIEW/current.png
controls@rossa:~ 0$ 

The "current.png" is a link to the most recent snapshot, and is what you get when you "viewsnap".

Below is an example medm snippet of what could be used in screens to enable this functionality.  I've fixed up a couple of the screens, but there are a lot more that need to be updated.

"shell command" {
    object {
        x=666
        y=597
        width=40
        height=40
    }
    command[0] {
        label="Settings and Procedures"
        name="emacs"
        args="./help/IFO_ALIGN.txt &"
    }
    command[1] {
        label="View Snapshot"
        name="/opt/rtcds/caltech/c1/medm/snap/viewsnap"
        args="MISC/IFO_ALIGN.adl &"
    }
    command[2] {
        label="View old snapshots"
        name="/opt/rtcds/caltech/c1/medm/snap/viewold"
        args="MISC/IFO_ALIGN.adl &"
    }
    command[3] {
        label="Update Snapshot"
        name="/opt/rtcds/caltech/c1/medm/snap/updatesnap"
        args="MISC/IFO_ALIGN.adl &"
    }
    clr=14
    bclr=30
}
  7061   Tue Jul 31 19:34:55 2012 KojiUpdateSTACISOpen loop gains and block diagram

With your definition of the open loop gain, G=+1 is the condition to have singularity in a closed loop transfer function 1/(1-G).

But this is not the sole criteria of the loop stability.
Basically, the closed loop transfer function should not have "unphysical" pole.
For more about loop instability, you should refer stability criteria in literature such as Nyquist's stability criterion.

Both of the X and Z loops look unstable with the current gain.

  7060   Tue Jul 31 18:59:59 2012 JamieUpdateCDSupdated medm paths for MISC (IFO,CDS,VIDEO), IFO alignment scripts updated accordingly

In an attempt to clean up the medm situation, I did a bunch of further rearrangement and cleanup.  Instead of having c1ifo, I moved a bunch of stuff into a MISC directory, including all of the CDS, IFO, and VIDEO screens:

controls@rossa:~ 0$ ls -1 /opt/rtcds/caltech/c1/medm/MISC | grep -v _BAK
CDS_BIO_STATUS.adl
CDS_FE_STATUS.adl
CDS_IPC_ERR.adl
help
ifoalign
IFO_ALIGN.adl
ifoalign.orig
IFO_CONFIGURE.adl
IFO_CONFIGURE.txt
IFO_OVERVIEW.adl
IFO_OVERVIEW_AIDAN.adl
IFO_STATE.adl
snap
VIDEO.adl
controls@rossa:~ 0$

I updated the sitemap and the relevant screens accordingly.

I also updated the IFO_ALIGN script infrastructure a bit.  The new IFO ALIGN location is /opt/rtcds/caltech/c1/medm/MISC/ifoalign.  The scripts are now called simply:

/opt/rtcds/caltech/c1/medm/MISC/ifoalign/misalign_soft.csh
/opt/rtcds/caltech/c1/medm/MISC/ifoalign/restore_soft.csh
/opt/rtcds/caltech/c1/medm/MISC/ifoalign/save_soft.csh

These run Jenne's new soft restore stuff, and store burt snapshots for optic alignment in /opt/rtcds/caltech/c1/medm/MISC/ifoalign/burt.

This has all been checked into the svn.

  7059   Tue Jul 31 15:33:17 2012 MashaConfigurationPEMGurlap Pin Map

I checked the connections specified in the old Gulap Pin Map and found that they do not correspond to the current values. I mapped out the current connections (in this case, the letter refers to the labeled pin on the mil/spec while the number refers to the pin on the 37 pin DSub, labeled consecutively):

A-1, B-2, C-3, D-4, E-5, F-6, G-7, H-Unused, J-8, K-unused, L-9, M-10, N -11, P-12, S-13, T-Unused, U-14, V-15, W-16, X-17, Y-18, Z-Unused, a-Unused, b-19, c-20, UnlabeledPin-Unused.

There are 20 pins in use of 26 total, which is good because that means Jenne and I can use the ~70m long 24 wire cable to make a new Gurlap 1 cable.

GurlapPinMap3.png

  7058   Tue Jul 31 15:24:53 2012 YaakovUpdateSTACISOpen loop gains and block diagram

First, a quick note on the PZT I thought I killed- it was most likely something in the high voltage amplifier that broke, since I put the amplifier in another STACIS with a working y-axis PZT and it still didn't work properly. Conclusion: something in the y-axis amplifier circuitry is broken, not the PZT itself.

Today I retook the open loop gains in the X and Z axes (Y axis out of commission for now, see above). With the loop open, I input a swept sine signal from 0.1 to 100 Hz, and measure the output of the geophones. This way all the transfer function that are present in the closed loop are present here as well: the transfer functions of the physical STACIS, the geophone pre-amplifier circuit, the high-voltage amplifier, and the PZT actuators.

Here is a block diagram showing what I am measuring, with the various transfer functions in blue boxes (the measurement is their product):

 stacis_block.bmp

x_OL.bmpz_OL.bmp

z_OL.fig

x_OL.fig

These open loop gains show there is gain of at least 10x from 2 to 80 Hz in the z-axis and 2 to 60 Hz in the x-axis. This is the region I was seeing isolation in when I switched to closed loop, which is consistent. These measurements were with all the pots in the geophone preamplifier set very low, so more gain (and thus isolation) is hypothetically possible if I find a way to stop the horizontal axes from becoming unstable at higher gains. There is unity gain at around 0.5 Hz and 100 Hz for the z-axis, but the phase is nowhere near 180 deg. at these points so there shouldn't be instability due to this. The peak at around 15 Hz is consistent with old records of the STACIS open loop gain.

  7057   Tue Jul 31 15:17:58 2012 JamieUpdateCDSc1ifo medm screens checked into CSD userapps svn

I moved the medm/c1ifo directory into the CDS userapps svn at cds/c1/medm/c1ifo, and then linked it back into the medm directory:

controls@rossa:~ 0$ ls -al /opt/rtcds/caltech/c1/medm/c1ifo
lrwxrwxrwx 1 controls controls 56 2012-07-31 11:53 /opt/rtcds/caltech/c1/medm/c1ifo -> /opt/rtcds/caltech/c1/userapps/release/cds/c1/medm/c1ifo
controls@rossa:~ 0$

I then committed whatever useful was in there.  We need to remember to commit when we make changes.

  7056   Tue Jul 31 08:01:22 2012 steveUpdateVACcc1 switched

Quote:

Quote:

Our cold cathode 423 gauges are 10 years old. They get insulated by age and show no ionization current.  We should replace them at the next vent. I'm buying 4 at $317ea

 

 We have two CC1 gauges at the pump spool. One is in vertical position and the spare is in horizontal. Today I switched the cable from vertical to horizontal CC1

This one is reading higher pressure. It is an old dirty gauge. It may trigger the RGA protection when it's reading goes above 1e-5 torr

The real pressure is 2e-6 torr

 cc1h spikes closed VM1 last night. VM1 is opened.

Attachment 1: cc1h1d.png
cc1h1d.png
  7055   Tue Jul 31 00:27:52 2012 DenUpdatePEMtrillium

We have a Trillium for several days from Vladimir. I've put seismometer inside the foam box on linolium. I was not able to level the seismometer on granite as this Trillium does not have level screws. Does anybody know where they are?  Readout box stands on the foam box as seismometer cable is short (~2 meters).

Cables go to STS1 inputs (7-9) on ADC 3.

  7054   Mon Jul 30 22:52:51 2012 YaakovUpdateSTACISGeophone calibration

Tonight I looked at the signal from a geophone and accelerometer side by side, in order to see if they show the same ground motion and if the sensitivity factor I am using to convert from V to m/s is right. This is plotted below, along with the current estimates for accelerometer and geophone noise:

sensor_comp.bmp

sensor_comp.fig

From this it is pretty clear that at least one of the sensitivity factors (V/m/s) I am using is wrong (the noise levels are much lower than the ground motion, so they can't account for the difference). I suspect it is the geophone one, because Wilcoxon provided these sensitivities for each individual accelerometer, but I was just using the number I found in online specs for the geophones.

The reason the online value is wrong is probably because of the value of the shunt resistor, a resistor that just goes across the top of the geophone (its purpose is to provide damping, by Lenz's Law). The specs assume a value of 380 Ohm, but I measured the one in the STACIS to be about 1.85 kOhm.

Assuming the accelerometer signal is correct, I multiplied the geophone signal by different factors to try to get an idea of what the true calibration factor is, and found that a value of 0.25 (m/s)/V gives decent agreement at higher frequencies (below 10 Hz the sensitivity drops off, according to the online specs). This is shown below:

sensor_comp2.bmp

sensor_comp2.fig

Above, the geophone noise was recalculated with the new sensitivity and assuming that both geophones in the noise measurement had the same sensitivity. I took the transfer function of two geophones side by side to see if their gains were dramatically different; this plot is shown below. The coherence is only good for a small band, but looking at that band the gain is approximately unity, implying very roughly that the sensitivity of each is approximately the same. The lack of coherence is strange, and I'm not sure what the cause is. Even using the shaker near the geophones only improved the coherence slightly.

2_geo_gain.bmp2_geo_coherence.bmp

Attachment 2: sensor_comp2.bmp
  7053   Mon Jul 30 17:24:45 2012 YaakovUpdateSTACISRevised sensor noise plot; dead PZT

The geophone noise in my last eLog was taken before any amplification of the signal, but what really matters is the noise after amplification, since it is this signal that the PZTs are driven by. The noise goes on to be amplified about 1000x before the geophone signal gets to the PZTs.

To obtain a more relevant noise plot, I multiplied the geophone noise by 1,000, the approximate gain of the amplification stage for the geophones (called the "compensator board", the semicircular board that sits toward the top of the STACIS). Below is a plot (sensor_noise.fig) that shows the noise for the geophones after amplification and the accelerometer noise (with accelerometers set with a gain of 100x, their highest).

sensor_noise.bmp

The actual signal from both these sensors has the right magnitude to drive the PZTs (whereas it was much too small in my last plot, where I looked at the sensors before any gain)- this means that for these sensors, both of which are outputting signals that are ready to provide feedback to the STACIS, the accelerometer noise is significantly lower than the geophone noise. This is good news, because it means that there could be a real advantage to using the accelerometers instead of the geophones.

In the process of investigating further advantages of the accelerometers, I believe I killed one of the horizontal PZTs in the spare STACIS (the eBay one). The story: I had that axis in closed loop, and I saw the STACIS shudder, heard a noise, and there was a faint acrid smell. I shut the STACIS off and took out the high voltage card at the base but couldn't find any visible signs of damage (like the current-limiting resistors which burn when a PZT shorts, acc. to old STACIS records). I then tried driving the PZTs with a sine wave, and there was no response in that axis (the other axes looked fine), which leads me to believe I either did unseen damage to the high voltage amplifier (for the y-axis) or killed the PZT itself.

Attachment 1: sensor_noise.bmp
  7052   Mon Jul 30 16:05:36 2012 DenUpdatedigital noisefilter checker

 We decided to write a script that will check online filters for digital noise. One method can be implemented using the following algorithm:

  • calculate filter output using single precision
  • calculate filter output using double precision and assume that it is precise
  • find digital noise at the output of the filter when single precision is used
  • extrapolate the result to the double precision filter dividing by 2D-S ~ 107, D - number of bits used in double precision mantissa, S - in single precision

Restriction: Single precision filter internal variables must be checked for overflows.

I applied this method to filtering a 1 Hz sine wave with a notch filter. Precise output should also be a 1 Hz wave => at other frequencies we see noise => digital noise spectrum should coincide with filter output. The plot shows the method worked out for this example.

iir_psd_notch.png

Using this method I estimated digital noise of butter("LowPass", 2, 0.001) applied to white noise. Sampling frequency was 16 kHz. 

iir_psd_lp.png iir_time_lp.png

  7051   Mon Jul 30 15:49:16 2012 steveUpdateVACcc1 switched

Quote:

Our cold cathode 423 gauges are 10 years old. They get insulated by age and show no ionization current.  We should replace them at the next vent. I'm buying 4 at $317ea

 

 We have two CC1 gauges at the pump spool. One is in vertical position and the spare is in horizontal. Today I switched the cable from vertical to horizontal CC1

This one is reading higher pressure. It is an old dirty gauge. It may trigger the RGA protection when it's reading goes above 1e-5 torr

The real pressure is 2e-6 torr

Attachment 1: cc1changedfVtH.png
cc1changedfVtH.png
  7050   Mon Jul 30 14:24:25 2012 DenUpdatedigital noisebiquad key is working

Quote:

What is "DQF"?  Is that the biquad?  And what is the difference between DF1 and DF2?  Why don't you just write out the name, so it's more clear.

DQF - biquad form
DF1 - direct form 1
DF2 - direct form 2
LNF - low-noise form
 
The difference between them is described in Matt's slides G0900928-v1. I think, LNF coefficients are incorrect in the presentation
 
lnf.png
  7049   Mon Jul 30 12:38:45 2012 JamieUpdateCDSMove to RCG 2.5 tag release

I moved the RCG to the advLigoRTS-2.5 tag:

controls@c1iscey ~ 0$ ls -al /opt/rtcds/rtscore/release
lrwxrwxrwx 1 controls 1001 19 Jul 30 12:02 /opt/rtcds/rtscore/release -> tags/advLigoRTS-2.5
controls@c1iscey ~ 0$ 

There are only very minor differences between what we were running on the the 2.5 branch.  I have NOT rebuilt all the models yet.

I was hoping that there was something in the tagged release that might fix this hard-crash-on-filter-load issue that we're seeing in the c1tst model.  It didn't.  Still have no idea what's going on there.

  7048   Mon Jul 30 10:05:29 2012 JenneUpdatePSLPMC is still not fixed

Quote:

The PMC was locking right the way, but it's transmission would not go up.  Finally I get it back up by moving the "sticky" DC Gain slider up and down a few times.

 The FSS was -2.9, and the PMC won't lock happily unless you bring this back to 0.  The symptom that this is happening is that the PMC reflection camera is totally saturated, but the PMC still looks like it's locked on 00.

  7047   Mon Jul 30 09:00:18 2012 steveUpdatePSLPMC is still not fixed

The PMC was locking right the way, but it's transmission would not go up.  Finally I get it back up by moving the "sticky" DC Gain slider up and down a few times.

Attachment 1: pmcA.png
pmcA.png
Attachment 2: pmcB.png
pmcB.png
  7046   Fri Jul 27 16:32:17 2012 JamieUpdateCDSRCG bug exposed by simple c1tst model

As Den mentioned in 7043, attempting to run the c1tst model was causing the entire c1iscey machine to crash.  Alex came over this morning and we spend a couple of hours trying to debug what was going on.

c1tst is the simplest possible model you can have: 1 ADC and 2 filter modules.  It compiles just fine, but when you tried to load it the machine would completely freeze.

We eventually tracked this down to a non-empty filter file for one of the filter modules.  It turns out that the model was crashing when it attempted to load the filter file.  Once we completely deleted all the filters in the module, the model would run.  But then if you added back a filter to the filter file and tried to "load coefficients", the model/machine would immediately crash again.

So it has something to do with the loading of the filter coefficients from the filter file.  We tried different filters and it didn't seem to make a difference.  Alex thought it might have something to do with zeros in some of the second-order sections, but that wasn't it either.

There's speculation that it might be related to a very similar bug that Joe reported at LLO a month ago: https://bugzilla.ligo-wa.caltech.edu/bugzilla/show_bug.cgi?id=398

Things we tried, none of which worked:

  • adding a DAC
  • turning on/off biquad
  • disabling the float denormalization fix

This is a real mystery.  Alex and I are continuing to investigate.

  7045   Fri Jul 27 14:30:49 2012 JamieUpdatedigital noisebiquad key is working

What is "DQF"?  Is that the biquad?  And what is the difference between DF1 and DF2?  Why don't you just write out the name, so it's more clear.

  7044   Fri Jul 27 14:28:36 2012 steveUpdateVACcold cathode gauges

Our cold cathode 423 gauges are 10 years old. They get insulated by age and show no ionization current.  We should replace them at the next vent. I'm buying 4 at $317ea

 

Attachment 1: cc_gauges.png
cc_gauges.png
  7043   Fri Jul 27 14:27:14 2012 JamieUpdateCDSnew c1tst model for testing RCG code

Quote:

 

 I wanted to test biquad form in this model. I added biquad=1 flag to cdsParameters, compiled, installed and restarted it. After that c1iscey suspended.

The same thing as we had several month ago

controls@c1iscey /opt/rtcds/caltech/c1/target/c1tst/c1tstepics 0$ cat iocC1.log

Starting iocInit
iocRun: All initialization complete
sh: iniChk.pl: command not found
Failed to load DAQ configuration file

I have fixed the iniChk.pl issue (which actually fixed a separate model startup-on-boot issue that we had been having).  However, that is completely unrelated to the system freeze.  I'll discuss that in a separate post.

  7042   Thu Jul 26 21:31:44 2012 DenUpdatedigital noiseonline biquad works

Quote:

   Still I do not like huge (2n+1) harmonics in the output of the biquad filter, I do not get them in the simulations. They are absent in the time series as well. So this is not a psd-estimation effect. 

 Excitation generator created these harmonics. When I applied low-pass butterworth filter, I've got the result of online filtering close to simulations. On the second graph blue is biquad filter output spectrum, red corresponds to DF2. 1 Hz sin wave was filtered with a notch filter of Q=100, depth=300 at 1 Hz.

df2_bqf_lp.png    df2_bqf_lp_spec.png

  7041   Thu Jul 26 17:39:49 2012 DenUpdatedigital noisebiquad key is working

 I've filtered a 1 Hz sin wave excitation with a notch filter inside c1sus and c1rfm models. The biquad key is switched on in the last one, c1sus uses DF2. The results are indeed different.

df2_bqf.png    df2_bqf_spec.png

Still I do not like huge (2n+1) harmonics in the output of the biquad filter, I do not get them in the simulations. They are absent in the time series as well. So this is not a psd-estimation effect.

iir_time_notch.pngiir_psd_notch.png

 

  7040   Thu Jul 26 16:08:59 2012 YaakovUpdateSTACISNoise plot update

I have a tentative noise plot for the STACIS that includes accelerometer noise, geophone noise, and platform motion with the STACIS off. (Accelerometer noise was measured for the VEL and NONE setting, which are settings on the accelerometer box which make the accelerometer signal correspond to velocity and acceleration, respectively. ) I'm focusing on sensor noise because this is the variable I am looking at changing, and knowing how the sensor noise translates into STACIS platform motion is therefore important.

stacy_noise.bmp

stacy_noise.fig

The accelerometer and geophone noise I determined as described in my last eLog (http://nodus.ligo.caltech.edu:8080/40m/7027) Along the way I found out several things of importance:

1) Horizontal geophones are ONLY horizontal geophones. This is obvious in retrospect, because the springs supporting the magnet inside must be oriented based on vertical/horizontal operation.

2) The geophones in the STACIS are GS-11D (geospace), with a sensitivity of 32 V/m/s (compared to about 3.9 V/m/s for the accelerometers in VEL setting).

3) The accelerometers have different V/m/s sensitivities. I noticed the voltage output of one was consistently higher than the other, leading to very high noise estimates, but then Jenne showed me the actual calibration factors of the individual accelerometers which differed by as much as 0.4 V/m/s (a few percent difference). Taking this into account made the noise plots much more reasonable, but variations in calibration could still create some error.

The accelerometer noise agrees fairly well with the specs on the Wilcoxon page (http://www.wilcoxon.com/prodpdf/731A-P31%2098079a1.pdf). The geophone noise seems surprisingly low; it is even better than the geophone below about 4 Hz. 

To see how this noise translates into actual platform motion, I took PSDs of the STACIS while it was off, on with accelerometer feedback, and on with geophone feedback (the "off" PSD is in the above noise plot). Using this data I'm working on estimating a transfer function that shows how the sensor noise translates to motion so I can come up with a sensor noise budget.

feedbacks.bmp

feedbacks.fig

This shows that the geophones are actually doing a better job of isolating than the accelerometers, which is not surprising if the noise plot is accurate and the geophones are actually lower noise. It must be noted, though, that the noise plot was for the horizontal geophones whereas the plot above is for the vertical axis which may have a different noise level. Also, the vertical have some extra isolation by being enclosed in a metal stack with rubber padding at its base.

The problem with the STACIS in the past was the differential motion it introduced. I think this might be because the horizontal isolation was not uniform for each chamber. This means that even what would be symmetric motion (no differential length change) would be translated to differential motion because one end is more fixed than the other. Having accelerometers or better-padded geophones (maybe like the vertical geophones) in the STACIS ought to help with this by making the horizontal isolation more consistent and thus reducing differential motion. So the key may not be the geophone noise as much as varying geophone sensitivities or variation in how well they're mounted in the STACIS. I can test this by swapping out the horizontal geophones with other spares, changing the tightness of the mount, and seeing if either of these changes the horizontal isolation significantly, since these are factors that may differ from unit to unit.

I will also compare horizontal closed loop response with geophone vs. accelerometer feedback to see if the geophones are only doing a good job in the above plot because of their extra padding (the vertical stack).

  7039   Thu Jul 26 15:43:03 2012 ranaUpdateLSCmodelled ringdown

You cannot use the digital system for this. You hook up a scope to the transmitted light as well as the incoming light (after the MC, perhaps at IP_POS). Then you acquire the data from both places simultaneously using an ethernet equipped scope. The step response of the PDs used for this has to be calibrated separately.

  7038   Thu Jul 26 13:10:51 2012 janoschUpdateLSCmodelled ringdown

We fitted shutter and ringdown functions to the ringdown data. It is not perfectly clear how the power change due to the shutter is handed over to the power change due to ringdown. The fit suggests that the ringdown starts at a later time, but this does not necessarily make sense. It could be that the slow power decrease when the shutter starts clipping the TEM00 beam is followed by the cavity, but then the power decrease becomes too fast when the shutter reaches the optical axis and the ringdown takes over. Also, the next measurement should be taken with adjusted DC offset.

Ringdown.png

  7037   Thu Jul 26 12:10:28 2012 DenUpdateCDSnew c1tst model for testing RCG code

Quote:

I made a new model, c1tst, that we can use for debugging the FREQUENT RCG bugs that we keep encountering.  It's a bare model that runs on c1iscey.  Don't do any thing important in here, and don't leave it in some crappy state.  Clean if up when you're done.

 I wanted to test biquad form in this model. I added biquad=1 flag to cdsParameters, compiled, installed and restarted it. After that c1iscey suspended.

The same thing as we had several month ago

controls@c1iscey /opt/rtcds/caltech/c1/target/c1tst/c1tstepics 0$ cat iocC1.log

Starting iocInit
iocRun: All initialization complete
sh: iniChk.pl: command not found
Failed to load DAQ configuration file

  7036   Thu Jul 26 10:22:03 2012 DenUpdatedigital noisenotch, lowpass filters

Quote:

If the problem is the precision in DTT, then why would the noise change when the corner frequency of the filter is changed?

And how about checking for filter noise in the situation we saw online? 4th order low pass at 1 Hz or 8 Hz.

 This is because when we plot signals with sampling frequencies 2k and 16k with the same BW, we actually create psd/coherence using different numbers of points in FFT calculations as NFFT ~ fs/bw, fs-sampling frequency. So we secretly used 8 times more fft points for 16k signal then for 2k signal. Following plots demonstrate this effect. The first plot shows transfer function and coherence for filtering of 16k signal with butter('LowPass',4,8) and 2k signal with butter('LowPass',4,1)  when BW=0.1. There is a disturbance in coherence for 2k signal below 2 Hz. Now let's take BW=0.8 and plot transfer function and coherence for 16k signal. We can see the same effect of coherence disturbance.

same_bw.png    16384_bw0p8.png

The similar effect takes place when we change the cut-off frequency. The following plots show transfer function and coherence of two pairs of 2kHz signals. 4 order butterworth low-pass filter was used. For the first pair cut-off frequency was 1 Hz, for the second 10 Hz.  On the first plot BW=0.1 and there is a disturbance in coherence below 1 Hz. However on the second plot when BW=0.01, this effect is lost.

corners_bw0p1.png     corners_bw0p01.png

I guess my goal is to figure out when these effects come from fft calculations and when from digital filter noise.

  7035   Thu Jul 26 02:44:17 2012 JenneUpdateIOOCentering the MCR camera

[Yaakov, Jenne]

The short version:

Rana and Koji pointed out to us that the MCR camera view was still not good.  There were 2 problems:

(1) Diagonal stripes through the beam spot.  Yuta and I saw this a week or 2 before he left, but we were bad and didn't elog it, and didn't investigate.  Bad grad students.

(2) Clipping of the left side of the beam (as seen on the monitors).  This wasn't noticed until Yaakov's earlier camera work, since the clipped part of the beam wasn't on the monitor.

The fix for #1 was to partially close the iris which is the first "optic" the beam sees on the AP table after leaving the vacuum. 

The "fix" for #2 was that the wrong beam has been going to the camera for an unknown length of time.  We picked the correct beam, and all is well again.

We moved the 10% BS that splits the main beam into the (MC REFL PD) path and the (MCR camera + WFS) path.  It looked like the transmission through there was close to the edge of the BS.  We didn't think that this was causing the clipping that we saw on the camera (since when we stepped MC1 in Yaw, the beam spot had to move a lot before we saw any clipping), but it seemed like a good idea to make the beam not near the edge of the optic, especially since, being a 2" optic, there was plenty of room, and we were only using ~half of the optic.  We didn't touch anything else in the WFS path, since that looks at the transmission through this BS, but we had to realign the beam onto MC REFL.

The long version:

(1)  The fix for #1 was to partially close the iris which is the first "optic" the beam sees on the AP table after leaving the vacuum.  It looks like that's why the iris was there in the first place.  When we found it this evening, the iris was totally open, so my current theory is that someone was on the AP table doing something, and accidentally bumped the handle for the iris, then left it completely open when they realized that they had touched it.  I think Steve had been doing something on the AP table around then, but since Yuta and I didn't elog our observation (bad grad students!), I can't correlate it with any of Steve's elogs. We were not able to find exactly where this "glow" that the iris is used to obscure comes from, but we traced it as far as the viewport, so there's something going on inside.

(2)  The "fix" for #2 was that the wrong beam has been going to the camera for an unknown length of time.  We picked the correct beam, and all is well again. 

We spent a long time trying to figure out what was going on here.  Eventually, we moved the camera around to different places (i.e. right before the MC REFL PD, with some ND filters, and then we used a window to pick off a piece of the beam right as it comes out of the vacuum before going through the iris, put some ND filters, then the camera).  We thought that right before the MC REFL PD was also being clipped, indicating that the clipping was happening in the vacuum (since the only common things between the MC REFL PD path and the camera path are the iris, which we had removed completely, and a 2" 10% beam splitter.  However, when we looked at a pickoff of the main beam before any beamsplitters, we didn't see any evidence of clipping.  I think that when we had the camera by MC REFL, we could have been clipping on the ND filters that I had placed there.  I didn't think to check that at the time, and it was kind of a pain to mush the camera into the space, so we didn't repeat that.  Then we went back to the nominal MCR camera place to look around.  We discovered that the Y1 which splits the camera path from the WFS path has a ghost beam which is clipping on the top right side (as you look at the face of the optic) of the optic, and this is the beam that was going to the camera (it's a Y1 since we only want a teensy bit of light to go to the camera, the rest goes to the WFS).  There is another beam which is the main beam, going through the center of the optic, which is the one which also reflects and heads to the WFS.  This is the beam which we should have on the camera.  Yaakov put the camera back in it's usual place, and put the correct beam onto the center of the camera.  We did a check to make sure that the main beam isn't clipping, and when I step MC1 yaw, the beam must move ~1.5mm before we start to see any clipping on the very edge of the beam.  To see / measure this, we removed the optic which directs the beam to the camera, and taped an IR card to the inside of the black box.  This is ~about the same distance as to the nominal camera position, which means that the beam would have to move by 1.5mm on the camera to see any clipping.  The MC REFL PD is even farther from MC1 than our IR card, so the beam has to fall off the PD before we see the clipping.  Thus, I'm not worried about any clipping for this main beam.  Moral of the story, if you made it this far:  There wasn't any clipping on any beams going to either the WFS or the MC REFL PD, only the beam going to the camera.

  7034   Wed Jul 25 23:03:22 2012 ranaUpdatedigital noisenotch, lowpass filters

If the problem is the precision in DTT, then why would the noise change when the corner frequency of the filter is changed?

And how about checking for filter noise in the situation we saw online? 4th order low pass at 1 Hz or 8 Hz.

  7033   Wed Jul 25 22:16:36 2012 KojiUpdateLSCringdown measurement

Is this the step response of the single pole low pass???
It should have an exponential decay, shouldn't it? So it should be easier to comprehend the result with a log scale for vertical axis...

I think you need a fast shutter. It is not necessary to be an actual shutter but you can use faster thing
which can shut the light. Like PMC or IMC actuators.

Another point is that you may like to have a witness channel like the MC transmission to subtract other effect.

  7032   Wed Jul 25 17:35:44 2012 LizUpdateComputer Scripts / ProgramsSummary Pages are in the right place!

The summary pages can now be accessed from the "Daily Summary" link under LOGS on the 40 meter Wiki page.

  7031   Wed Jul 25 16:55:01 2012 DenUpdatedigital noisenotch, lowpass filters

 Direct Form 2 is noisy in the first test. This is the one similar to Matt's in his presentation. Input signal was a sine wave at 1 Hz with small amplitude white noise x[n] = sin(2*pi*1*t[n]) + 1e-10 * random( [-1, 1] ). It was filtered with a notch filter: f=1Hz, Q=100, depth=210dB. SOS representation was calculated in Foton. Sampling frequency is 16kHz.

iir_psd.png        iir_time.png   iir_coh.png

DF2 output noise level is the same if I change white noise amplitude while DF1, BQF, LNF can follow it. Time series show quantization noise of DF2. I've plotted coherence of the signals relative to DF1, because non of the signals will be coherent to it at low frequencies due to fft calculations.  

In the second test the input was white noise  x[n] = random( [-1, 1] ) It was filtered with a 2 order low-pass butterworth filter with cut-off frequency f = 0.25 Hz. SOS representation was calculated in Python. Sampling frequency is 16kHz.

iir_psd_lowpass.png         iir_time_lowpass.png      iir_coh_lowpass2.png

In this test all implementations work fine. I guess dtt works with single precision and for that reason we see disturbance in coherence when we do the same test online.

  7030   Wed Jul 25 16:31:01 2012 JenneUpdateLSCYarm green locking to arm - PDH box saturating

Quote:

... it was not possible to get the cavity locked in green. So we decided to do the first measurement with infrared locked only. 

 When we sat down to align the Yarm to the green, the green light was happy to flash in the cavity, but wouldn't lock, even after Jan had tweaked the mirrors such that we were flashing the TEM00 mode.  When we went down to the end to investigate, the Universal PDH box was saturating both negative and positive.  Turning down the gain knob all the way to zero didn't change anything, so I put it back to 52.5.  Curiously, when we unplugged the Servo OUT monitor cable (which was presumably going to the rack to be acquired), the saturation happened much less frequently.  I think (but I need to look at the PDH box schematic) that that's just a monitor, so I don't know why that had to do with anything, but it was repeatable - plug cable in, almost constant saturation....unplug cable, almost no saturation.

Also, even with the cable unplugged, the light wouldn't flash in the cavity.  When I blocked the beam going to the green REFL PD (used for the PDH signal), the light would flash.

Moral of the story - I'm confused.  I'm going to look up the PDH box schematic before going back down there to investigate.

  7029   Wed Jul 25 15:33:55 2012 janoschUpdateLSCringdown measurement

We did our first ringdown measurement on the Y arm. First we tried to keep the arm locked in green during the ringdown, but for some reason it was not possible to get the cavity locked in green. So we decided to do the first measurement with infrared locked only.

For the measurement we had to change the LSC model to acquire the C1:LSC-TRY_OUT_DQ at higher sampling frequency. We changed the sampling frequencies of C1:LSC-{TRX,TRY}_OUT_DQ from 2048Hz to 16384Hz.

The measurement was done at GPS 1027289507. The ringdown curve looks very clean, but there seem to be two time constants involved. The first half of the curve is influenced by the shutter speed, then curvature is changing sign and the ringdown is likely taking over. We will try to fit a curve to the ringdown part, but it would certainly be better to have a faster shutter and record a more complete ringdown.

Rindown1.png

 

ELOG V3.1.3-